### v1

# GraphLoRA: Structure-Aware Contrastive Low-Rank Adaptation for Cross-Graph Transfer Learning

Abstract.
Graph Neural Networks (GNNs) have demonstrated remarkable proficiency in handling a range of graph analytical tasks across various domains, such as e-commerce and social networks. Despite their versatility, GNNs face significant challenges in transferability, limiting their utility in real-world applications. Existing research in GNN transfer learning overlooks discrepancies in distribution among various graph datasets, facing challenges when transferring across different distributions. How to effectively adopt a well-trained GNN to new graphs with varying feature and structural distributions remains an under-explored problem. Taking inspiration from the success of Low-Rank Adaptation (LoRA) in adapting large language models to various domains, we propose GraphLoRA, an effective and parameter-efficient method for transferring well-trained GNNs to diverse graph domains. Specifically, we first propose a Structure-aware Maximum Mean Discrepancy (SMMD) to align divergent node feature distributions across source and target graphs. Moreover, we introduce low-rank adaptation by injecting a small trainable GNN alongside the pre-trained one, effectively bridging structural distribution gaps while mitigating the catastrophic forgetting. Additionally, a structure-aware regularization objective is proposed to enhance the adaptability of the pre-trained GNN to target graph with scarce supervision labels. Extensive experiments on six real-world datasets demonstrate the effectiveness of GraphLoRA against eleven baselines by tuning only $20\%$ of parameters, even across disparate graph domains. The code is available at https://anonymous.4open.science/r/GraphLoRA.
graph neural network, low-rank adaptation, transfer learning
* Corresponding author.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Graph Neural Networks (GNNs) have emerged as a powerful tool for processing and analyzing graph-structured data, demonstrating exceptional performance across diverse domains (e.g., e-commerce (Zhou et al., 2021), social networks (Wang et al., 2018), and recommendation (Yang et al., 2022; Lai et al., 2024)) for diverse tasks, such as node classification (Chen et al., 2020b), link prediction (Zhang et al., 2020), and graph classification (Zhou et al., 2020).
Figure 1. Negative transfer occurs in cross-graph adaption, where PubMed, CiteSeer, and Cora are citation networks, whereas Photo and Computer are co-purchase networks.
Despite their superiority in capturing intricate graph relationships, GNNs face significant challenges when confronting substantial variations in underlying data distributions (Cao et al., 2023; Liu et al., 2022). Transferring a well-trained GNN to another graph typically yields suboptimal performance (i.e., negative transfer (Zhao et al., 2024)) due to out-of-distribution issues. As depicted in Figure 1, negative transfer occurs in cross-graph adaption, even when transferring between citation networks or co-purchase networks within a same domain.
To tackle this challenge, a common approach is to imbue GNNs with generalizable graph knowledge through the ’pre-train and fine-tune’ paradigm (Qiu et al., 2020; Han et al., 2021). In this approach, it’s crucial to integrate domain-specific knowledge while preserving the universal knowledge acquired during pre-training. Consequently, Parameter Efficient Fine-Tuning (PEFT) has garnered considerable attention for its ability to mitigate the risks of overfitting and catastrophic forgetting (Xu et al., 2023). PEFT updates only a small portion of the parameters while keeping the remaining parameters frozen, thus mitigating the risk of forgetting the universal knowledge (Xu et al., 2023).
For instance, Gui et al. (Gui et al., 2024) propose a structure-aware PEFT method named G-Adapter, aimed at adapting pre-trained Graph Transformer Networks to various graph-based downstream tasks. Li et al. (Li et al., 2024) propose the AdapterGNN method, which applies the adapter to non-transformer GNN architectures. However, while these methods focus on incorporating PEFT into GNNs, they lack specific mechanisms to address discrepancies in distribution among different graphs, such as variations in node features and graph structures. As a result, they encounter challenges when applied to graphs with varying distributions.
How to effectively adapt well-trained GNNs to graphs with different distributions remains an under-explored problem, posing a non-trivial task due to three major challenges. (1) Cross-graph feature discrepancy. The node feature distributions between source and target graphs can vary significantly, impeding the transferability of pre-trained GNNs. For example, attributes in academic citation networks (e.g., PubMed (Yang et al., 2016)) differ greatly from those in e-commerce co-purchase networks (e.g., Computer (Shchur et al., 2018)). (2) Cross-graph structural discrepancy. The structural characteristics of source and target graphs are often diverged. For instance, academic citation networks typically exhibit higher density and consists of more cyclic motifs compared to e-commerce networks (Wu et al., 2008; Srivastava, 2010). (3) Target graph label scarcity. The effectiveness of pre-trained GNNs often relies on sufficient labels on target graphs, which are not always available in the real-world. For instance, in social networks, typically only a small fraction of high-degree nodes are labeled (Wang et al., 2024).
To this end, in this paper, we present GraphLoRA, a structure-aware low-rank contrastive adaptation method for effective transfer learning of GNNs in cross-graph scenarios. Specifically, we first introduce a Structure-aware Maximum Mean Discrepancy (SMMD) to minimize the feature distribution discrepancy between source and target graphs. Moreover, inspired by the success of Low-Rank Adaptation (LoRA) (Hu et al., 2022) in adapting large language models to diverse natural language processing tasks and domains (Long et al., 2023; Bao et al., 2023), we construct a small trainable GNN alongside the pre-trained one coupled with a tailor-designed graph contrastive loss to mitigate structural discrepancies. Additionally, we develop a structure-aware regularization objective by harnessing local graph homophily to enhance the model adaptability with scarce labels in the target graph.
The main contributions of this work are summarized as follows:
  • We propose a novel strategy for measuring feature distribution discrepancy in graph data, which incorporates the graph structure into the measurement process.
  • We propose GraphLoRA, a novel method tailored for cross-graph transfer learning. The low-rank adaptation network, coupled with graph contrastive learning, efficiently incorporates structural information from the target graph, mitigating catastrophic forgetting and addressing structural discrepancies across graphs. Furthermore, we theoretically demonstrate that GraphLoRA possesses robust representational capabilities, facilitating effective cross-graph transfer.
  • We propose a structure-aware regularization objective to enhance the adaptability of pre-trained GNN to target graphs, particularly in contexts with limited label availability.
  • Extensive experiments on real-world datasets demonstrate the effectiveness of our proposed method by tuning a small fraction of parameters, even cross disparate graph domains.
2. Related Work
2.1. Graph Transfer Learning
Graph transfer learning involves pre-training a GNN and applying it to diverse tasks or datasets. Common techniques in graph transfer learning include multi-task learning (Hwang et al., 2020; Han et al., 2021), multi-network learning (Jiang, 2021; Ni et al., 2018), domain adaptation (Dai et al., 2023; Liu et al., 2023a), and pre-train fine-tune approaches (Qiu et al., 2020; Zhang et al., 2022a; Li et al., 2024; Zhu et al., 2023). However, multi-task learning, multi-network learning and domain adaptation are typically employed for cross-task transfer or necessitate direct relationships between the source and target graphs, which is not applicable to our problem (Han et al., 2021; Ni et al., 2018; Liu et al., 2023a). Therefore, we focus on pre-train and fine-tune techniques, involving pre-training a GNN on the source graph and subsequently fine-tuning it on the target graph. For instance, GCC (Qiu et al., 2020), GCOPE (Zhao et al., 2024), and GraphFM (Lachi et al., 2024) focus on pre-training to develop more general GNNs. In contrast, DGAT (Guo et al., 2024) is dedicated to designing architectures that perform better in out-of-distribution scenarios. GTOT (Zhang et al., 2022a), AdapterGNN (Li et al., 2024), and GraphControl (Zhu et al., 2023) focus on fine-tuning, aiming to adapt pre-trained GNNs to various graphs. Most relevant to our work is GraphControl, which freezes the pre-trained GNN and utilizes information from the target graph as a condition to fine-tune the newly added ControlNet for cross-domain transfer. In contrast to GraphControl, our method aligns the node feature distributions to facilitate the transfer of the pre-trained GNN, rather than using node attributes as conditions. Additionally, we leverage graph contrastive learning to facilitate knowledge transfer and utilize graph structure knowledge to enhance the adaptability of the pre-trained GNN.
Figure 2. The framework of GraphLoRA: Fine-tuning the pre-trained GNN for the target graph. The node feature adaptation and structural knowledge transfer learning modules are designed to alleviate feature and structural discrepancies, respectively. Furthermore, the structure-aware regularization objective is crafted to enhance the adaptability of the pre-trained GNN.
2.2. Parameter-Effcient Fine-Tuning
”Pre-train and fine-tune” has emerged as the predominant paradigm in transfer learning (Sun et al., 2023). Despite its success, full fine-tuning is frequently inefficient and susceptible to challenges like overfitting and catastrophic forgetting (Gui et al., 2024; Li et al., 2024; Xu et al., 2023). In recent years, PEFT has emerged as an alternative, effectively mitigating these issues while achieving comparable performance (Xu et al., 2023). PEFT methods update only a small portion of parameters while keeping the remaining parameters frozen. For instance, Adapter tuning (Houlsby et al., 2019; Lin et al., 2020) inserts trainable adapter modules into the model, while Prompt tuning (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021b) inserts trainable parameters into the input or hidden states of the model. BitFit (Zaken et al., 2022) updates only the bias terms in the model, while LoRA (Hu et al., 2022) applies low-rank decomposition to reduce the number of trainable parameters. Recently, some efforts have been made to introduce PEFT into GNNs. For example, methods like GPPT (Sun et al., 2022), GPF (Fang et al., 2023), GraphPrompt (Liu et al., 2023b), and ProG (Sun et al., 2023) utilize prompt tuning to adapt pre-trained GNNs across diverse downstream tasks. G-Adapter (Gui et al., 2024) adapts pre-trained Graph Transformer Networks for various graph-based downstream tasks, whereas AdapterGNN (Li et al., 2024) applies adapters to non-transformer GNN architectures.
3. Preliminaries
3.1. Notations and Background
In this paper, we utilize the notation $\mathcal{G}=\left(\mathcal{V},\mathcal{E}\right)$ to denote a graph, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{N}\}$ represents the node set with $N$ nodes in the graph, and =$\mathcal{E}=\{(v_{i},v_{j})|v_{i},v_{j}\in\mathcal{V}\}$ represents the edge set in the graph. The feature matrix is denoted as $\mathbf{X}\in\mathbb{R}^{N\times d}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the node feature of $v_{i}$, and $d$ is the dimension of features. Furthermore, the adjacency matrix of the graph is denoted as $\mathbf{A}\in\{0,1\}^{N\times N}$, where $\mathbf{A}_{i,j}=1$ iff $(v_{i},v_{j})\in\mathcal{E}$. To avoid confusion, we employ the superscripts $s$ and $t$ in this paper to distinguish between the source and target graphs.
Graph neural networks. A major category of GNNs is message-passing neural networks (MPNNs) (Gilmer et al., 2017). MPNNs follow the ”propagate-transform” paradigm, which can be described as follows:
(1) $\bar{\boldsymbol{h}}_{s}^{l}$ $=\mathrm{Propagate}_{l}\left(\left\{\boldsymbol{h}_{t}^{l-1}|v_{t}\in\mathcal{N}\left(v_{s}\right)\right\}\right),$
(2) $\boldsymbol{h}_{s}^{l}$ $=\mathrm{Transform}_{l}\left(\boldsymbol{h}_{s}^{l-1},\bar{\boldsymbol{h}}_{s}^{l}\right),$
where $\mathcal{N}\left(v_{s}\right)$ denotes the neighboring node set of node $v_{s}$, $\boldsymbol{h}_{s}^{l}$ represents the node embedding of node $v_{s}$ in the $l$-th layer, and $\bar{\boldsymbol{h}}_{s}^{l}$ denotes the aggregated representation from neighboring nodes.
Low-Rank Adaptation (LoRA). LoRA (Hu et al., 2022) is a widely used PEFT methods, designed to efficiently fine-tune LLMs across tasks and domains. Compared to the Adapter method, LoRA provides better performance without introducing additional inference latency (Hu et al., 2022). Specifically, for a pre-trained weight matrix $\boldsymbol{W}_{0}\in\mathbb{R}^{m\times n}$, its weight update is expressed through a low-rank decomposition, given by $\boldsymbol{W}_{0}+\bigtriangleup\boldsymbol{W}=\boldsymbol{W}_{0}+\boldsymbol{BA}$, where $\boldsymbol{B}\in\mathbb{R}^{m\times r}$, $\boldsymbol{A}\in\mathbb{R}^{r\times n}$, and the rank $r\ll\min\left(m,n\right)$. During fine-tuning, the pre-trained weight matrix $\boldsymbol{W}_{0}$ is frozen, while $\boldsymbol{B}$ and $\boldsymbol{A}$ are tunable. The low-rank adaptation strategy effectively reduces the number of parameters requiring fine-tuning while maintaining high model quality without introducing inference latency. Notably, by sharing the vast majority of model parameters, it enables quick task switching and allows the pre-trained model to be applied to various tasks and domains.
3.2. Problem Statement
Given a GNN $g_{\theta}$ pre-trained on the source graph $\mathcal{G}^{s}$, our goal is to obtain an optimal GNN $f_{\varTheta}^{\star}$ for the target graph $\mathcal{G}^{t}$,
(3) $f_{\varTheta}^{\star}=\underset{\varTheta}{\text{argmin}}\,\,\mathcal{L}\left(f_{\varTheta}\left(\mathbf{X}^{t},\mathbf{A}^{t}\right),Y_{train}\right),$
where $Y_{train}$ denotes training labels, and $\mathcal{L}$ is the model tuning loss function. $f_{\varTheta}\left(\cdot\right)=h_{\varTheta}\circ g_{\theta}\left(\cdot\right)$, with $h_{\varTheta}$ as the tunable module and $g_{\theta}$ frozen. We focus on the node classification task in this work.
4. Methodology
4.1. Framework Overview
The overall framework of the model is illustrated in Figure 2. Firstly, a node feature adaptation module is designed to perform feature mapping on the target graph. Within this module, we propose a Structure-aware Maximum Mean Discrepancy strategy to minimize discrepancy in node feature distributions between the source and target graphs. After that, we introduce a structural knowledge transfer learning module to mitigate structural disparity. Taking inspiration from LoRA, we apply low-rank adaptation to the pre-trained GNN. This can be seen as incorporating an additional GNN $g_{\phi}^{{}^{\prime}}$ with the same architecture, but utilizing the weight updates as its parameters. Additionally, we utilize graph contrastive learning to facilitate knowledge transfer. During this process, we freeze the weights of $g_{\theta}$ and fine-tune $g_{\phi}^{{}^{\prime}}$. To enhance the adaptability of the pre-trained GNN to scenarios with scarce labels, a structure-aware regularization objective is proposed to effectively leverage the structural information of the target graph. Finally, we employ multi-task learning to optimize multiple objectives.
4.2. Node Feature Adaptation
Previous works on transfer learning have suggested that minimizing the discrepancy in feature distributions between the source and target datasets is crucial for effective knowledge transfer (Ben-David et al., 2010; Zhuang et al., 2021). To achieve this, a projector is designed to perform feature mapping on the node features of the target graph, as follows:
(4) $\boldsymbol{z}_{i}^{t}=p\left(\boldsymbol{x}_{i}^{t};\boldsymbol{\omega}\right),$
where $p\left(\cdot;\boldsymbol{\omega}\right)$ is the projector with parameters $\boldsymbol{\omega}$, and $\boldsymbol{z}_{i}^{t}\in\mathbb{R}^{d^{s}}$.
To optimize the projector, our goal is to minimize the discrepancy in feature distributions between $\boldsymbol{x}^{s}$ and $\boldsymbol{z}^{t}$. In the realm of domain adaptation, a commonly employed metric to quantify the dissimilarity between two probability distributions is the Maximum Mean Discrepancy (MMD) (Borgwardt et al., 2006; Gretton et al., 2012). The fundamental principle underlying MMD is to measure the dissimilarity between two probability distributions by comparing their means in a high-dimensional feature space. Specifically, it can be expressed as follows:
(5) $\begin{split}\mathcal{L}_{mmd}=&\small{\frac{1}{\left(N^{t}\right)^{2}}\sum_{i=1}^{N^{t}}{\sum_{i^{\prime}=1}^{N^{t}}{k\left(\boldsymbol{z}_{i}^{t},\boldsymbol{z}_{i^{\prime}}^{t}\right)}}}-\small{\frac{2}{N^{t}N^{s}}\sum_{i=1}^{N^{t}}{\sum_{j=1}^{N^{s}}{k\left(\boldsymbol{z}_{i}^{t},\boldsymbol{x}_{j}^{s}\right)}}}\\ +&\small{\frac{1}{\left(N^{s}\right)^{2}}\sum_{j=1}^{N^{s}}{\sum_{j^{\prime}=1}^{N^{s}}{k\left(\boldsymbol{x}_{j}^{s},\boldsymbol{x}_{j^{\prime}}^{s}\right)}}},\end{split}$
where $k(\cdot,\cdot)$ denotes the kernel function.
For the optimization of $\mathcal{L}_{mmd}$, we can observe that the first term of $\mathcal{L}_{mmd}$ maximizes the distance between node features in the target graph, while the second term minimizes the distance between node features of the source and target graphs. The third term denotes a constant. However, the node features in graph data are not independently and identically distributed ($i.i.d.$), $i.e.$, exhibiting correlation with the graph structure. For instance, neighboring nodes tend to exhibit similar features, which is overlooked by $\mathcal{L}_{mmd}$. Therefore, it is crucial to consider the graph structure when minimizing the discrepancy in feature distributions between $\boldsymbol{x}^{s}$ and $\boldsymbol{z}^{t}$. This aids in retaining the structural information in node features, such as homophily, during feature mapping.
Specifically, we introduce Structure-aware Maximum Mean Discrepancy (SMMD), which incorporates graph structure into the measurement of distribution discrepancy. In particular, smaller weights are assigned to node pairs with stronger connections, and larger weights are assigned to node pairs with weaker connections.
First, it is crucial to quantify the strength of relationships between node pairs. Since the adjacency matrix only provides a local perspective on the graph structure (Hassani and Ahmadi, 2020), we utilize the graph diffusion technique to transform the adjacency matrix into a diffusion matrix (Klicpera et al., 2019; Li et al., 2023a). The diffusion matrix allows us to evaluate the strength of relationships between node pairs from a global perspective, facilitating the discovery of potential connections between node pairs and preserving them during feature mapping. Specifically, the diffusion matrix is defined as:
(6) $\boldsymbol{S}=\sum_{r=0}^{\infty}{\psi_{r}\boldsymbol{T}^{r}}.$
where $\boldsymbol{T}$ represents the transition matrix, which is related to the adjacency matrix $\boldsymbol{A}^{t}$, and $\psi_{r}$ represents the weight coefficient. We utilize a popular variant of the diffusion matrix, Personalized PageRank (PPR) (Page et al., 1998), which employs $\boldsymbol{T}=\boldsymbol{A}^{t}\boldsymbol{D}^{-1}$ and $\psi_{r}=\alpha\left(1-\alpha\right)^{r}$, where $\boldsymbol{D}$ denotes the diagonal degree matrix, i.e. $\boldsymbol{D}_{i,i}=\sum_{j=1}^{N^{t}}{\boldsymbol{A}_{i,j}^{t}}$, and $\alpha\in\left(0,1\right)$ represents the teleport probability. The elements $\boldsymbol{S}_{i,j}$ in the obtained diffusion matrix $\boldsymbol{S}$ indicate the strength of relationships between node $v_{i}^{t}$ and node $v_{j}^{t}$.
Finally, we define the Structure-aware Maximum Mean Discrepancy loss function as follows:
(7) $\mathcal{L}_{smmd}=$ $\frac{1}{\sum_{i=1}^{N^{t}}\sum_{i^{\prime}=1}^{N^{t}}\gamma_{i,i^{\prime}}}\sum_{i=1}^{N^{t}}\sum_{i^{\prime}=1}^{N^{t}}\gamma_{i,i^{\prime}}k\left(\boldsymbol{z}_{i}^{t},\boldsymbol{z}_{i^{\prime}}^{t}\right)$
$-\frac{2}{N^{t}N^{s}}\sum_{i=1}^{N^{t}}\sum_{j=1}^{N^{s}}k\left(\boldsymbol{z}_{i}^{t},\boldsymbol{x}_{j}^{s}\right)+\frac{1}{(N^{s})^{2}}\sum_{j=1}^{N^{s}}\sum_{j^{\prime}=1}^{N^{s}}k\left(\boldsymbol{x}_{j}^{s},\boldsymbol{x}_{j^{\prime}}^{s}\right),$
(8) $\gamma_{i,i^{\prime}}=$ $\log\left(1+\frac{1}{\boldsymbol{S}_{i,i^{\prime}}}\right),$
where $\mathcal{L}_{smmd}$ incorporates graph structure during computation.
4.3. Structural Knowledge Transfer Learning
Recent study (Cao et al., 2023) suggests that the disparity in graph structure between the source and target graphs impedes the transferability of pre-trained GNNs. Straightforward approaches such as full parameter fine-tuning of pre-trained GNNs may also lead to additional issues such as catastrophic forgetting (Pfeiffer et al., 2021). Consequently, effectively transferring the pre-trained GNN to target graphs becomes a formidable challenge when there is a significant discrepancy in graph structure.
Drawing inspiration from the success of LoRA across various tasks and domains, we propose incorporating adaptation for pre-trained weights, as depicted in Figure 2. During fine-tuning, we freeze the pre-trained weights while allowing newly added parameters to be tunable. From another perspective, it can be seen as maintaining the network architecture and parameters of the pre-trained GNN while introducing an additional GNN with the same architecture and utilizing the weight updates as its parameters.
Formally, let $g_{\phi}^{{}^{\prime}}$ represents the newly added GNN, $GNN_{l}\left(\cdot;\boldsymbol{W}^{l}\right)$ and $GNN_{l}^{{}^{\prime}}\left(\cdot;\bigtriangleup\boldsymbol{W}^{l}\right)$ denote the $l$-th layer of $g_{\theta}$ and $g_{\phi}^{{}^{\prime}}$, respectively, where $\boldsymbol{W}^{l},\bigtriangleup\boldsymbol{W}^{l}\in\mathbb{R}^{d^{l-1}\times d^{l}}$ are parameter matrices. The output of GNN at the $l$-th layer is modified from $\boldsymbol{H}^{l}=GNN_{l}\left(\boldsymbol{H}^{l-1};\boldsymbol{W}^{l}\right)$ to $\boldsymbol{H}^{l}=GNN_{l}\left(\boldsymbol{H}^{l-1};\boldsymbol{W}^{l}\right)+GNN_{l}^{{}^{\prime}}\left(\boldsymbol{H}^{l-1};\bigtriangleup\boldsymbol{W}^{l}\right)$, where $\boldsymbol{H}^{0}=\boldsymbol{Z}^{t}$, and $\boldsymbol{Z}^{t}$ represents the feature-mapped node feature matrix. Let $\boldsymbol{H}=g_{\theta}\left(\boldsymbol{Z}^{t}\right)$ and $\boldsymbol{H}^{{}^{\prime}}=g_{\phi}^{{}^{\prime}}\left(\boldsymbol{Z}^{t}\right)$ represent the output of $g_{\theta}$ and $g_{\phi}^{{}^{\prime}}$, respectively. We apply low-rank decomposition to the weight update to reduce the number of tunable parameters. Specifically, the output of the $l$-th layer of GNN is represented as follows:
(9) $\boldsymbol{H}^{l}=GNN_{l}\left(\boldsymbol{H}^{l-1};\boldsymbol{W}^{l}\right)+GNN_{l}^{{}^{\prime}}\left(\boldsymbol{H}^{l-1};\boldsymbol{W}_{B}^{l}\boldsymbol{W}_{A}^{l}\right),$
where $\boldsymbol{W}_{B}^{l}\in\mathbb{R}^{d^{l-1}\times r}$, $\boldsymbol{W}_{A}^{l}\in\mathbb{R}^{r\times d^{l}}$, and the rank $r\ll\min\left(d^{l-1},d^{l}\right)$.
The advantages of the above design are two-fold. First, the pre-trained GNN preserves general structural knowledge from the source graph, while the newly added one serves to incorporate specific structural information from the target graph, jointly facilitating downstream tasks. Keeping the pre-trained weights frozen during fine-tuning also mitigates the issue of catastrophic forgetting. Second, since the target graph may suffer from label scarcity, the low-rank decomposition reduces the number of tunable parameters to update and thus mitigates potential overfitting issues.
Unlike vanilla LoRA, which fine-tunes the network based on downstream task objectives, we further introduce graph contrastive learning to facilitate structural knowledge transfer. Specifically, we consider the embeddings obtained by two GNNs (frozen and tunable ones) for the same node as positive samples, while treating the embeddings for different nodes as negative samples. Furthermore, to enhance the learning effectiveness of node embeddings, we incorporate label information into graph contrastive learning. This involves treating the embeddings of nodes belonging to the same category as positive samples and considering the embeddings of nodes from different categories as negative samples.
Formally, the graph contrastive learning loss (van den Oord et al., 2018) is defined as
(10) $\mathcal{L}_{cl}=-\sum_{i=1}^{N^{t}}{\sum_{y_{i}=y_{k}}{\log\frac{e^{\rho\left(\boldsymbol{h}_{i},\boldsymbol{h}_{i}^{{}^{\prime}}\right)/\tau}+\varepsilon e^{\rho\left(\boldsymbol{h}_{i},\boldsymbol{h}_{k}^{{}^{\prime}}\right)/\tau}}{e^{\rho\left(\boldsymbol{h}_{i},\boldsymbol{h}_{i}^{{}^{\prime}}\right)/\tau}+\sum_{j\neq i}{e^{\rho\left(\boldsymbol{h}_{i},\boldsymbol{h}_{j}^{{}^{\prime}}\right)/\tau}}+\sum_{j\neq i}{e^{\rho\left(\boldsymbol{h}_{i},\boldsymbol{h}_{j}\right)/\tau}}}}},$
where $y_{i}$ is the category of node $v_{i}^{t}$, and $\varepsilon\in(0,1)$ is the weight. The low-rank adaptation network coupled with tailor-designed graph contrastive learning incorporates structural information from the target graph, maximizing the mutual information between the pre-trained GNN and the newly added GNN. Therefore, such a strategy mitigates structural discrepancies across graphs, facilitating the adaptation of pre-trained GNNs to target graphs.
We further provide theoretical justification for the robust representation capability of pre-trained GNNs with low-rank adaptation.
Theorem 1.
Let $\overline{g}$ be a target GNN with $\overline{L}$ layers and $g_{0}$ be an arbitrary frozen GNN with $L$ layers, where $\overline{L}\leqslant L$. Under mild conditions on ranks and network architectures, there exist low-rank adaptations such that the low-rank adapted model $g_{0}$ becomes exactly equal to $\overline{g}$.
The proof of Theorem 1, along with additional theoretical analysis, are provided in Appendix A. The theorem suggests that effective cross-graph transfer, i.e., $g^{\star}$ achieves an optimal solution, can be accomplished through low-rank adaptation applied to the pre-trained GNN $g_{0}$, thereby equating $g_{0}$ with $g^{\star}$.
4.4. Structure-aware Regularization
In real-world scenarios, the homophily phenomenon is prevalent in graph data, such as citation networks or co-purchase networks (Li et al., 2023b). In general, homophily reflects the tendency for ”like to attract like” (McPherson et al., 2001), indicating that connected nodes are prone to sharing similar labels (Li et al., 2023b; Shchur et al., 2018). In the cross-graph transfer learning context, we leverage the homophily principle to alleviate label scarcity in the target graph.
Inspired by previous work (Mao et al., 2021), we propose a structure-aware regularization objective based on the homophily principle of graph data. Specifically, we assume that the predicted label vectors of connected neighbors on the target graph are similar, while those of disconnected neighbors are dissimilar. In contrast to GraphSage (Hamilton et al., 2017), we utilize direct connected neighbors instead of random walk to better satisfy the assumption, which can be formulated as:
(11) $\begin{split}\mathcal{L}_{str}=&\sum_{i\neq j}[{\boldsymbol{A}_{i,j}^{t}\log\varsigma\left(sim\left(\hat{\boldsymbol{y}}_{i},\hat{\boldsymbol{y}}_{j}\right)\right)}\\ +&\left(1-\boldsymbol{A}_{i,j}^{t}\right)\log\left(1-\varsigma\left(sim\left(\hat{\boldsymbol{y}}_{i},\hat{\boldsymbol{y}}_{j}\right)\right)\right)],\end{split}$
where $\hat{\boldsymbol{y}}_{i}$ represents the predicted label vector of node $v_{i}^{t}$, $sim(\cdot,\cdot)$ represents the inner product, and $\varsigma(\cdot)$ represents the sigmoid function. Despite the limited availability of labeled data, the above regularization objective effectively utilizes the inherent homophily property in the graph to mitigate the challenge of label scarcity.
4.5. Optimization
Finally, we employ the following output layer to classify the target nodes based on the output of the GNN,
(12) $\tilde{y}_{i}=$ $c\left(\boldsymbol{h}_{i}+\boldsymbol{h}_{i}^{{}^{\prime}}\right),$
where $c(\cdot)$ represents the classifier, and $\tilde{y}_{i}=argmax\left(\hat{\boldsymbol{y}}_{i}\right)$ denotes the predicted class of node $v_{i}^{t}$. The classification loss function is defined as follows:
(13) $\mathcal{L}_{cls}=$ $-\frac{1}{N^{t}}\sum_{i}{y_{i}\log\tilde{y}_{i}+\left(1-y_{i}\right)\log\left(1-\tilde{y}_{i}\right)}.$
After acquiring the pre-trained GNN from the source graph, we proceed to fine-tune it by utilizing the labeled data available on the target graph. To achieve this, we employ multitask learning to jointly optimize multiple objective functions. The overall objective function is defined as follows:
(14) $\mathcal{L}=\mathcal{L}_{cls}+\lambda_{1}\mathcal{L}_{smmd}+\lambda_{2}\mathcal{L}_{cl}+\lambda_{3}\mathcal{L}_{str}+\lambda_{4}\left\|\varTheta\right\|,$
where the last term acts as a regularization term to prevent overfitting, and the weight coefficients $\lambda_{1-4}$ determine the importance of each objective function in the overall optimization process.
Table 1. Comparison of experimental results in public and 10-shot settings. The notations ”-PM,” ”-CS,” ”-C,” ”-P,” and ”-Com” represent the pre-training datasets PubMed, CiteSeer, Cora, Photo, and Computer, respectively. The best experimental results are highlighted in bold, while the second-best results are underscored with a underline.
Method PubMed CiteSeer Cora Photo Computer
public 10-shot public 10-shot public 10-shot public 10-shot public 10-shot
GCN 78.66±0.56 73.28±0.93 70.50±0.75 64.52±1.40 82.00±0.97 71.88±1.26 92.17±0.75 85.39±1.57 86.66±1.20 71.97±1.16
GAT 78.30±0.43 74.96±0.67 70.94±1.08 64.14±2.12 80.58±1.50 72.04±0.61 92.91±0.22 86.38±0.65 86.80±0.71 74.82±2.36
COSTA 79.94±1.16 76.98±1.29 70.36±1.29 65.56±1.94 81.84±0.92 76.28±1.42 92.04±0.58 83.02±0.43 87.00±0.33 71.28±0.82
CCA-SSG 80.58±0.85 78.76±1.62 73.76±0.75 67.46±0.92 83.94±1.02 76.84±0.80 92.74±0.33 85.56±0.67 88.08±0.35 76.94±1.37
HomoGCL 81.04±0.05 79.42±0.04 71.38±0.04 65.40±0.00 82.40±0.00 75.70±0.00 92.43±0.08 84.30±0.35 87.75±0.29 76.85±0.41
GPPT 77.78±0.31 74.84±0.55 67.56±0.33 64.14±0.63 80.16±0.38 72.94±0.24 92.10±0.23 86.32±0.56 88.34±0.18 77.30±0.44
GRACE 79.52±0.16 75.86±0.11 70.34±0.21 67.70±0.00 82.28±0.04 76.40±0.00 92.32±0.31 86.16±0.02 85.54±0.30 74.39±0.03
GRACEt-PM 79.44±0.15 75.80±0.16 67.74±0.05 57.40±0.00 76.82±0.24 64.44±0.30 92.14±0.14 85.95±0.10 84.81±0.41 76.15±0.09
GRACEt-P 71.10±0.00 57.70±0.62 58.00±0.00 49.38±0.04 72.40±0.00 57.88±0.04 92.25±0.34 86.20±0.02 84.24±0.21 74.57±0.01
GTOT-PM 76.48±1.12 71.92±0.66 69.96±1.31 60.10±0.54 78.82±1.07 68.54±0.44 90.18±0.88 83.69±0.52 84.88±0.31 64.60±3.28
GTOT-P 76.44±0.63 71.42±0.70 69.28±1.09 60.86±0.80 79.40±2.26 69.04±0.75 90.42±0.42 83.37±0.97 84.66±0.42 64.60±1.19
AdapterGNN-PM 76.44±0.97 72.78±0.72 62.76±1.42 58.64±0.44 75.54±1.46 63.82±1.44 92.39±0.32 88.24±0.39 88.00±0.18 75.54±0.65
AdapterGNN-P 72.94±0.42 63.44±0.84 64.20±0.43 53.02±0.86 75.50±1.33 57.64±2.76 92.58±0.50 88.18±0.29 87.62±0.38 75.00±0.52
GraphControl-PM 78.30±0.43 75.96±1.00 69.02±1.65 60.82±0.41 77.84±0.67 69.32±2.11 90.73±0.75 86.65±0.51 85.94±0.96 74.47±2.43
GraphControl-P 74.58±1.99 58.94±0.69 59.12±1.34 53.36±1.92 73.46±1.73 65.72±0.44 90.67±0.50 86.23±0.59 86.11±0.50 71.86±2.38
GraphLoRA-PM 80.86±0.39 78.06±0.59 74.20±0.47 74.62±0.57 82.42±0.40 78.08±0.3 93.00±0.36 88.34±0.51 87.70±0.63 76.54±0.39
GraphLoRA-CS 80.64±0.43 78.08±0.67 74.08±0.26 74.80±0.6 82.00±0.23 78.30±0.46 93.08±0.11 89.00±0.43 87.72±0.45 76.44±0.05
GraphLoRA-C 80.38±0.50 77.78±0.55 73.98±0.45 74.62±0.65 82.00±0.80 78.00±0.51 92.92±0.29 88.69±0.35 88.10±0.31 76.45±0.48
GraphLoRA-P 78.46±0.67 74.84±1.67 72.80±0.58 72.02±1.64 81.28±0.40 76.54±0.54 92.47±0.33 88.89±0.89 87.35±0.49 75.27±0.22
GraphLoRA-Com 79.02±0.77 74.44±1.36 72.72±0.35 72.12±1.29 80.84±1.05 76.44±0.9 92.42±0.22 89.07±0.22 87.28±0.31 75.35±1.26
4.6. Complexity Analysis
In this section, we analyze the time complexity of GraphLoRA. For a target graph with $N^{t}$ nodes and $M$ edges, the node feature adaptation module performs feature mapping with a runtime of $\mathcal{O}(N^{t})$. By leveraging fast approximations (Andersen et al., 2006; Wei et al., 2018), the diffusion matrix $\boldsymbol{S}$ can be obtained in $\mathcal{O}(N^{t})$. $\mathcal{L}_{smmd}$ necessitates calculating distances between node pairs in each batch, achievable in $\mathcal{O}(N^{t}b)$ through the utilization of sampling techniques, where $b$ denotes the batch size. Similarly, in the structural knowledge transfer learning module, $\mathcal{L}_{cl}$ requires calculating similarity between node pairs in each batch, also with a complexity of $\mathcal{O}(N^{t}b)$. As for the structure-aware regularization objective, $\mathcal{L}_{str}$ considers connected nodes as positive samples and samples a small subset of nodes as negative samples, resulting in a complexity of $\mathcal{O}(M)$. In summary, the fine-tuning time complexity of GraphLoRA is $\mathcal{O}(N^{t}b+M)$, which is lightweight considering that the batch size is typically small.
5. Experiments
In this section, we conduct extensive experiments on benchmark datasets to evaluate GraphLoRA’s effectiveness in cross-graph transfer learning, aiming to answer the following research questions:
  • RQ1:
    How effective and efficient is GraphLoRA?
  • RQ2:
    Is GraphLoRA sensitive to hyperparameters?
  • RQ3:
    How do different modules contribute to its effectiveness?
  • RQ4:
    Can GraphLoRA mitigate catastrophic forgetting?
  • RQ5:
    Can GraphLoRA learn more distinguishable node representations than other baselines?
5.1. Experimental Setup
5.1.1. Datasets
We evaluate GraphLoRA on six publicly available datasets: three citation networks (PubMed, CiteSeer, and Cora) (Yang et al., 2016), two co-purchase networks (Photo and Computer) (Shchur et al., 2018), and one large-scale network (Reddit) (Hamilton et al., 2017). The statistics and detailed descriptions of these datasets are provided in Appendix B.
5.1.2. Baselines
We compare GraphLoRA with various baselines, including supervised methods GCN (Kipf and Welling, 2017) and GAT (Velickovic et al., 2018), graph contrastive learning methods GRACE (Zhu et al., 2020), COSTA (Zhang et al., 2022b), CCA-SSG (Zhang et al., 2021), and HomoGCL (Li et al., 2023b), as well as graph prompt learning method GPPT (Sun et al., 2022), and transfer learning methods GRACEt, GTOT (Zhang et al., 2022a), AdapterGNN (Li et al., 2024) and GraphControl (Zhu et al., 2023). Among these, GRACEt involves pretraining a GNN on the source graph using GRACE and then transferring it to the target graph for testing. Detailed descriptions of the baselines are provided in Appendix C.
5.1.3. Settings
For GraphLoRA, we use a standard two-layer GAT model as the backbone. The projector $p\left(\cdot;\boldsymbol{\omega}\right)$ and classifier $c(\cdot)$ are implemented with a single linear layer. We pre-train the GNN using GRACE and fine-tune it on the target graph using our method. Experiments are conducted in the public setting with sufficient labels, and in the 5-shot and 10-shot settings with limited labels. In the public setting, we partition citation networks using public splits (Yang et al., 2016), where each category has only 20 available labels for training. For co-purchase networks and Reddit, we randomly split the dataset into training (10%), validation (10%), and testing (80%) sets. In the 5-shot and 10-shot settings, each category in the training set contains only 5 and 10 labels, respectively. For both scenarios, 80% of the dataset is allocated for testing, while the remaining data is used as the validation set. For all methods, we conduct the experiments five times and report the average classification accuracy and standard deviation. The settings for baselines and additional results are provided in the Appendix due to space constraints.
5.2. Performance Comparison (RQ 1)
The performance of GraphLoRA on node classification tasks in both the public and 10-shot settings is presented in Table 1. Due to space constraints, we present only the experimental results of transfer learning methods on the pre-training datasets PubMed and Photo. The comprehensive table can be found in Appendix E.1, Table 7. GraphLoRA achieves either the best or second-best performance in most cases, underscoring its effectiveness. Compared to non-transfer learning scenarios, transfer learning scenarios are more challenging. Nevertheless, GraphLoRA achieves an average improvement of 1.01% over the best baseline results and 3.33% over GRACE. Specifically, it achieves an average improvement of 2.23% over GRACE in the public setting and 4.43% in the 10-shot setting. GraphLoRA shows a more significant performance improvement in the 10-shot setting, underscoring its effectiveness in scenarios with scarce labels.
5.2.1. Cross-graph Transfer Learning
From Table 1, we can observe that transfer learning methods exhibit poorer performance compared to non-transfer learning methods, highlighting the significant challenge of cross-graph transfer. In contrast, GraphLoRA demonstrates impressive transfer learning capabilities, even in cross-domain scenarios. Specifically, GraphLoRA achieves an average improvement of 10.12% over GRACEt, indicating that direct transfer of pre-trained GNNs results in suboptimal performance. Additionally, GraphLoRA achieves average improvements of 8.21% over GTOT, 9.78% over AdapterGNN, and 8.74% over GraphControl.
(a) Average accuracy across different shots.
(b) GraphLoRA improvements over GRACE and GRACEt.
Figure 3. Experimental results across different shots.
Table 2. Comparison of runtimes of different methods in public and 10-shot settings. For transfer learning methods, we report the fine-tuning runtimes.
Method PubMed CiteSeer Cora Photo Computer Avg.
public 10-shot public 10-shot public 10-shot public 10-shot public 10-shot
GCN 2.1s 1.7s 2.3s 2.3s 2.1s 2.1s 11.6s 11.2s 12.6s 12.1s 6.0s
GAT 2.8s 2.8s 2.8s 9.0s 2.5s 8.0s 14.3s 5.5s 16.1s 4.4s 6.8s
COSTA 1074.5s 524.0s 61.7s 127.4s 288.1s 135.2s 109.2s 71.8s 28.0s 77.1s 249.7s
CCA-SSG 8.7s 5.9s 6.8s 7.8s 4.5s 5.0s 4.9s 6.9s 13.4s 9.6s 7.3s
HomoGCL 166.9s 155.2s 6.6s 7.1s 5.1s 5.3s 26.8s 17.9s 100.8s 92.9s 58.5s
GPPT 23.6s 7.9s 36.5s 32.8s 37.5s 27.8s 181.7s 96.5s 378.7s 179.7s 100.3s
GRACE 88.5s 88.7s 6.4s 19.2s 16.9s 18.8s 31.7s 24.8s 64.6s 62.2s 42.2s
GRACEt 0.6s 0.4s 11.5s 12.0s 0.6s 12.0s 12.1s 4.1s 12.9s 5.8s 7.2s
GTOT 17.8s 79.3s 9.8s 17.6s 6.2s 10.9s 23.2s 19.9s 79.3s 56.3s 32.0s
AadpterGNN 18.9s 7.8s 14.6s 17.7s 11.6s 17.1s 45.3s 38.6s 106.9s 65.4s 34.4s
GraphControl 0.9s 1.1s 0.3s 1.5s 1.0s 0.8s 15.6s 9.8s 29.5s 13.0s 7.4s
GraphLoRA 43.7s 11.2s 5.0s 8.8s 10.7s 3.6s 44.6s 17.5s 108.9s 56.3s 31.0s
5.2.2. Scarce Labeling Impact on Performance
To further explore the impact of label scarcity on performance, we investigate the performance of GraphLoRA across the 1-shot to 10-shot setting, as illustrated in Figure 3. The figure reveals that, overall, GraphLoRA demonstrates a greater performance improvement compared to GRACE and GRACEt in scenarios with scarce labels. This observation not only reaffirms our earlier analysis but also substantiates the crucial role of transfer learning in scenarios with scarce labels. Furthermore, it is noteworthy that GraphLoRA consistently exhibits a more substantial performance improvement compared to GRACEt, providing additional confirmation that the direct transfer of pre-trained GNNs will result in suboptimal performance.
5.3. Efficiency Comparison (RQ 1)
Efficiency is a critical consideration in practical applications (Liu et al., 2021a). To evaluate the efficiency of GraphLoRA, we measure the runtime of different methods in both public and 10-shot settings on the same device, as depicted in Table 2. For transfer learning methods, we present the total runtime until model convergence during fine-tuning, while for other methods, we present the total runtime until model convergence during training. From Table 2, it is shown that the average runtime of GraphLoRA is lower than that of most baselines, indicating its high efficiency. It is noteworthy that GraphLoRA exhibits higher efficiency in the 10-shot setting compared to other baselines. This may be attributed to the effective mitigation of label sparsity through the structure-aware regularization objective, thereby facilitating easier model convergence.
Table 3. Experimental results on the Reddit dataset.
Method PM→R CS→R C→R P→R C→R R→R
GTOT 93.04±0.15 92.99±0.11 93.15±0.10 93.11±0.08 93.10±0.12 93.18±0.07
AdapterGNN 91.21±0.10 91.61±0.07 91.30±0.06 91.07±0.23 91.18±0.08 93.89±0.12
GraphControl 92.79±0.12 93.01±0.10 92.93±0.10 92.76±0.11 92.67±0.13 93.14±0.11
GraphLoRA 93.25±0.07 93.22±0.10 93.44±0.09 93.48±0.10 93.44±0.08 93.58±0.07
Table 4. Ablation experiment results in public and 10-shot settings. The best experimental results are highlighted in bold.
Variants PubMed CiteSeer Cora Photo Computer
public 10-shot public 10-shot public 10-shot public 10-shot public 10-shot
w/ mmd 79.46±0.82 76.44±1.76 73.12±0.43 74.08±0.91 81.46±0.54 76.80±0.86 92.70±0.37 87.91±1.03 87.75±0.12 76.01±0.24
w/o smmd 77.58±0.19 75.42±0.58 71.52±0.72 69.24±2.71 81.20±0.30 76.38±0.76 92.56±0.33 87.90±1.32 87.59±0.17 74.96±0.68
w/o cl 79.76±0.47 77.70±0.94 73.70±0.66 74.42±0.56 81.40±0.70 77.24±0.60 92.64±0.14 87.81±0.36 87.67±0.44 76.10±0.23
w/o str 79.88±0.49 77.54±0.83 71.12±0.55 68.12±0.42 80.64±0.65 74.50±1.34 92.63±0.49 87.96±0.83 87.60±0.27 75.87±0.67
w/o lrd 80.06±0.90 77.20±1.29 72.52±1.65 72.94±0.88 80.88±0.36 77.16±1.87 92.63±0.20 88.15±0.16 87.79±0.34 76.24±0.52
w/o nfa 79.72±0.13 76.86±0.09 68.20±0.25 27.76±0.65 75.24±0.55 69.68±0.24 87.26±0.82 83.17±0.34 80.97±0.82 69.46±1.06
w/o sktl 80.02±0.41 77.22±0.78 72.84±0.61 73.72±0.33 81.72±1.20 77.08±0.74 92.61±0.42 88.17±0.97 87.92±0.23 76.23±0.48
GraphLoRA 80.86±0.39 78.06±0.59 74.20±0.47 74.62±0.57 82.42±0.40 78.08±0.30 93.00±0.36 88.34±0.51 87.70±0.63 76.54±0.39
Figure 4. Performance across varying hyperparameter values.
5.4. Results on Large-Scale Dataset (RQ 1)
GraphLoRA can be easily applied to large-scale graphs using sampling techniques. We compare it with other transfer learning methods on the large-scale Reddit dataset, using the same sampling technique for all methods. The results are presented in Table 3, indicating that GraphLoRA performs well on large-scale graphs. Specifically, GraphLoRA outperforms all other methods in transfer learning scenarios and is second only to AdapterGNN in the non-transfer learning scenario. Although AdapterGNN excels in non-transfer learning scenarios, it performs poorly in transfer learning scenarios. Overall, GraphLoRA demonstrates superior performance, highlighting its effectiveness on large-scale graphs.
5.5. Hyperparameter Analysis (RQ 2)
5.5.1. Impact of $\lambda$
The model’s performance varies with different combinations of coefficients in the objective function. To investigate GraphLoRA’s sensitivity to hyperparameters, we conduct a parameter analysis on these coefficients. In our experiments, we tune the values of $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ within the range of [0.1,10]. The experimental results are presented in Figure 4, demonstrating that GraphLoRA’s performance remains generally stable, indicating low sensitivity to hyperparameters. Additionally, the impact of parameter adjustments on performance varies across datasets. For instance, the performance of GraphLoRA improves with an increase in $\lambda_{3}$ on the CiteSeer dataset. Conversely, its performance decreases with an increase in $\lambda_{3}$ on the PubMed and Computer datasets. To achieve optimal performance, strategies such as grid search, random search, and Bayesian optimization can be employed to obtain the best hyperparameter combination (Yu and Zhu, 2020). Overall, GraphLoRA demonstrates low sensitivity to hyperparameters, although the optimal parameter combination varies across different datasets.
5.5.2. Impact of $r$
The hyperparameter $r$ determines the parameter size of GraphLoRA. We evaluate GraphLoRA’s performance across $r$ values ranging from $2^{0}$ to $2^{8}$ in the public setting. The experimental results are depicted in Figure 4. The figure reveals that GraphLoRA maintains stable performance across different values of $r$. Even when $r$ is set to 1, GraphLoRA exhibits commendable performance, aligning with the understanding that a small $r$ value is adequate for LoRA (Hu et al., 2022). Notably, GraphLoRA experiences a decline in performance when $r$ is too small or too large. Generally, optimal performance is achieved when $r$ falls within the range of $2^{3}$ to $2^{5}$, with the tunable parameter ranging from 7% to 20%. This can be attributed to a small $r$ limiting parameters for effective fine-tuning, whereas a large $r$ may lead to overfitting due to an abundance of tunable parameters.
5.6. Ablation Studies (RQ 3)
To evaluate the effectiveness of each module in GraphLoRA, we compare it with seven different model variants. Specifically, ”w/ mmd” represents the method using the target term $\mathcal{L}_{mmd}$ rather than $\mathcal{L}_{smmd}$. ”w/o smmd”, ”w/o cl”, and ”w/o str” represent methods without using the target terms $\mathcal{L}_{smmd}$, $\mathcal{L}_{cl}$, and $\mathcal{L}_{str}$, respectively. Additionally, ”w/o lrd” is the method without employing low-rank decomposition for weight updates, while ”w/o nfa” and ”w/o sktl” represent methods without utilizing the node feature adaptation module and structural knowledge transfer learning module, respectively. The experimental results conducted in the public setting and 10-shot setting, following pre-training on the PubMed dataset, are depicted in Table 4. The results in the 5-shot setting can be found in Appendix E.3.
As illustrated in Table 4, it is shown that GraphLoRA consistently outperforms seven variants in most cases, thereby demonstrating the effectiveness of each module of GraphLoRA. Specifically, the most significant performance decline is observed for ”w/o nfa” and ”w/o smmd,” emphasizing the importance of considering the discrepancy in feature distributions in transfer learning. This observation further validates the effectiveness of our proposed Structure-aware Maximum Mean Discrepancy for measuring the discrepancy in node feature distributions. Moreover, ”w/o str” exhibits a more significant performance decline in the 10-shot setting compared to the public setting, indicating that the structure-aware regularization indeed contributes to improving the adaptability of pre-trained GNNs in scenarios with scarce labels.
5.7. Catastrophic Forgetting Analysis (RQ 4)
Table 5. Catastrophic forgetting analysis. After pre-training on the PubMed dataset, we fine-tune the model on other datasets and then test it back on the PubMed dataset.
[…]
FT 79.52±0.16 71.80±1.32 75.34±1.34 64.90±1.54 59.58±2.61
GTOT 76.48±1.12 78.28±0.33 72.28±0.94 67.82±2.50 61.58±2.07
AdapterGNN 76.44±0.97 77.50±0.68 76.50±0.73 74.06±1.95 73.34±1.56
GraphControl 78.30±0.43 78.00±0.27 75.16±2.09 65.48±1.02 65.42±3.29
GraphLoRA 80.86±0.39 79.84±0.28 79.82±0.24 80.06±0.28 79.88±0.28
(a) GRACE
(b) GRACEt
(c) GraphLoRA
Figure 5. Visualization of node embeddings on CiteSeer.
Fine-tuning the pretrained model with full parameters often leads to in catastrophic forgetting. To mitigate this, we freeze the pretrained parameters and introduce additional tunable parameters. To evaluate GraphLoRA’s ability to alleviate catastrophic forgetting, we first pre-train the model on the PubMed dataset, then fine-tune it on other datasets, and finally assess its performance back on PubMed. Experimental results comparing GraphLoRA with full parameter fine-tuning (FT) and other baselines are presented in Table 5. FT exhibits a significant performance decline, whereas GraphLoRA shows only a marginal decrease. Other baselines also experience performance declines, while less severe than that of FT. This is attributed to the fact that these methods freeze the pre-trained parameters while introducing additional trainable parameters, thus mitigating the issue of catastrophic forgetting to some degree. Moreover, GraphLoRA significantly outperforms FT (average 18.64%), highlighting its effectiveness in mitigating catastrophic forgetting.
5.8. Visualization of Representations (RQ 5)
In addition to quantitative analysis, we employ the t-SNE (Van der Maaten and Hinton, 2008) method to visually assess the GraphLoRA’s performance by visualizing the learned node embeddings on the CiteSeer dataset in the 10-shot setting. Specifically, Figure 5(a) shows embeddings learned by GRACE, while Figure 5(b) and Figure 5(c) display embeddings learned by GRACEt and GraphLoRA, respectively, following pre-training on the PubMed dataset. Each point in these figures represents a node, with its color denoting its label. From Figure 5, we observe that, compared to GRACE, the embeddings learned by GRACEt exhibit more blurred class boundaries, whereas the embeddings learned by GraphLoRA present clearer class boundaries. This observation suggests that GraphLoRA has a stronger capacity for learning node embeddings, proving beneficial for downstream tasks.
6. Conclusion
In this paper, we investigated the challenging problem of cross-graph transfer in graph neural networks. Drawing inspiration from the successful application of LoRA in fine-tuning large language models for diverse tasks and domains, we proposed a cross-graph transfer learning framework, GraphLoRA, for effective and parameter-efficient fine-tuning of pre-trained GNNs. We addressed the discrepancies in node feature distribution and graph structure between the source and target graphs by incorporating the node feature adaptation and structural knowledge transfer learning modules. Furthermore, we proposed a structure-aware regularization objective to enhance the adaptability of pre-trained GNNs for scenarios with scarce labels. Theoretical analysis guarantees that the GraphLoRA method possesses powerful representation capabilities and can precisely fit any target GNN under mild conditions. Extensive experiments on six real-world datasets demonstrate the effectiveness of GraphLoRA, even across disparate graph domains.
Appendix A Theoretical Analysis
A.1. Notations
We define $[N]\coloneqq\{1,2,\dots,N\}$. The SVD decomposition of matrix $\boldsymbol{W}$ is given as $\boldsymbol{W}=\boldsymbol{UDV}^{T}$, and $\sigma_{i}(\boldsymbol{W})=D_{i,i}$. The best rank-$r$ approximation (in the Frobenius norm or the 2-norm) of $\boldsymbol{W}$ is $\sum_{i=1}^{r}\sigma_{i}(\boldsymbol{W})\boldsymbol{u}_{i}\boldsymbol{v}_{i}^{T}$, where $\boldsymbol{u}_{i}$ and $\boldsymbol{v}_{i}$ are the $i$-th column of $\boldsymbol{U}$ and $\boldsymbol{V}$, respectively (Eckart and Young, 1936; MIRSKY, 1960; Zeng and Lee, 2023). Following (Zeng and Lee, 2023), we define the best rank-$r$ approximation as $LR_{r}(\boldsymbol{W})$. When $r\geqslant rank(\boldsymbol{W})$, it is obvious that $LR_{r}(\boldsymbol{W})=\boldsymbol{W}$. Considering the differences in architectures among various GNNs, for the sake of analytical simplicity, we consider the following general GNN architecture:
(15) $\mathrm{Propagate}:\bar{\boldsymbol{H}}^{l}=$ $\boldsymbol{H}^{l-1}\boldsymbol{P},$
(16) $\mathrm{Transform}:\boldsymbol{H}^{l}=$ $\mathrm{Re}LU(\boldsymbol{W}^{l}\bar{\boldsymbol{H}}^{l}+\boldsymbol{B}^{l}),$
where $\boldsymbol{B}^{l}=\boldsymbol{b}^{l}\mathbf{1},\mathbf{1}=[1,1,\dots,1]\in\mathbb{R}^{1\times N}$, and $\boldsymbol{P}$ represents the message transformation matrix associated with the adjacency matrix. For analytical simplicity, we assume that $(\boldsymbol{W}^{l})_{l=1}^{L}\in\mathbb{R}^{D\times D}$ and $(\boldsymbol{b}^{l})_{l=1}^{L}\in\mathbb{R}^{D\times 1}$.
We define an $L$-layer width-$D$ graph neural network as follows: $GNN_{L,D}(\cdot;(\boldsymbol{W}^{l})_{l=1}^{L},(\boldsymbol{b}^{l})_{l=1}^{L})\coloneqq\mathrm{ReLU}(\boldsymbol{W}^{L}\mathrm{ReLU}(\boldsymbol{W}^{L-1}\mathrm{ReLU}(\ldots)\boldsymbol{P}+\boldsymbol{B}^{L-1})\boldsymbol{P}+\boldsymbol{B}^{L})$. The target GNN $\overline{g}$, frozen GNN $g_{0}$, and adapted GNN $g$ are defined as follows:
(17) $\overline{g}=$ $GNN_{\overline{L},D}(\cdot;(\overline{\boldsymbol{W}}^{l})_{l=1}^{\overline{L}},(\overline{\boldsymbol{b}}^{l})_{l=1}^{\overline{L}}),$
(18) $g_{0}=$ $GNN_{L,D}(\cdot;(\boldsymbol{W}^{l})_{l=1}^{L},(\boldsymbol{b}^{l})_{l=1}^{L}),$
(19) $g=$ $GNN_{L,D}(\cdot;(\boldsymbol{W}^{l}+\bigtriangleup\boldsymbol{W}^{l})_{l=1}^{L},(\hat{\boldsymbol{b}}^{l})_{l=1}^{L}),$
where $\overline{L}\leqslant L$.
We define an $L$-layer width-$D$ fully connected neural network (FNN) as follows: $FNN_{L,D}(\cdot;(\boldsymbol{W}_{l})_{l=1}^{L},(\boldsymbol{b}_{l})_{l=1}^{L})\coloneqq\\ \mathrm{ReLU}(\boldsymbol{W}_{L}\mathrm{ReLU}(\boldsymbol{W}_{L-1}\mathrm{ReLU}(\ldots)+\boldsymbol{b}_{L-1})+\boldsymbol{b}_{L})$, where $(\boldsymbol{W}_{l})_{l=1}^{L}\in\mathbb{R}^{D\times D}$ are weight matrices, and $(\boldsymbol{b}_{l})_{l=1}^{L}\in\mathbb{R}^{D}$ are bias vectors. The target FNN $\overline{f}$, frozen FNN $f_{0}$, and adapted FNN $f$ are defined as follows:
(20) $\overline{f}=$ $FNN_{\overline{L},D}(\cdot;(\overline{\boldsymbol{W}}_{l})_{l=1}^{\overline{L}},(\overline{\boldsymbol{b}}_{l})_{l=1}^{\overline{L}}),$
(21) $f_{0}=$ $FNN_{L,D}(\cdot;(\boldsymbol{W}_{l})_{l=1}^{L},(\boldsymbol{b}_{l})_{l=1}^{L}),$
(22) $f=$ $FNN_{L,D}(\cdot;(\boldsymbol{W}_{l}+\bigtriangleup\boldsymbol{W}_{l})_{l=1}^{L},(\hat{\boldsymbol{b}}_{l})_{l=1}^{L}),$
where $\overline{L}\leqslant L$.
In addition, we define the partition $\mathcal{P}=\{P_{1},\cdots,P_{\overline{L}}\}=\{\{1,\cdots,\\ M\},\{M+1,\cdots,2M\},\cdots,\{(\overline{L}-1)M+1,\cdots,L\}\}$, where $M=\lfloor L/\overline{L}\rfloor$.
A.2. Expressive Power of Fully Connected Neural Networks with LoRA
Before presenting the theoretical analysis of the expressive power of graph neural networks with LoRA, we introduce the relevant lemmas from (Zeng and Lee, 2023) that discuss the expressive power of fully connected neural networks with LoRA.
Assumption 1.
For a fixed rank $R\in[D]$, the weight matrices of the frozen model $(\boldsymbol{W}_{l})_{l=1}^{L}$ and matrices $(\prod_{l\in P_{i}}{\boldsymbol{W}_{l}})+LR_{r}(\overline{\boldsymbol{W}}_{i}-\prod_{l\in P_{i}}{\boldsymbol{W}_{l}})$ are non-singular for all $r\leqslant R(M-1)$ and $i\in[\overline{L}]$.
Lemma 0.
Let $(\overline{\boldsymbol{W}}_{i})_{i=1}^{\overline{L}},(\boldsymbol{W}_{l})_{l=1}^{L}\in\mathbb{R}^{D\times D}$ matrices whose elements are drawn independently from arbitrary continuous distributions. Then, with probability 1, Assumption 1 holds $\forall R\in[D]$.
Lemma 0.
Under Assumption 1, if rank $R\geqslant\lceil\max_{i\in[\overline{L}]}rank(\overline{\boldsymbol{W}_{i}}-\prod_{l\in P_{i}}{\boldsymbol{W}_{l}})/M\rceil$, then there exists rank-$R$ or lower matrices $\bigtriangleup\boldsymbol{W}_{1},\cdots,\\ \bigtriangleup\boldsymbol{W}_{L}\in\mathbb{R}^{D\times D}$ and bias vectors $\hat{\boldsymbol{b}}_{1},\cdots,\hat{\boldsymbol{b}}_{L}\in\mathbb{R}^{D}$ such that the low-rank adapted model f can exactly approximate the target model $\overline{f}$, i.e., $f(\boldsymbol{x})=\overline{f}(\boldsymbol{x})$, $\forall\boldsymbol{x}\in\mathcal{X}$, where $\mathcal{X}$ is the input space.
Lemma 0.
Define $E_{i}=\sigma_{RM+1}(\overline{\boldsymbol{W}_{i}}-\prod_{l\in P_{i}}{\boldsymbol{W}_{l}})$, and $\xi=\\ \max(\max_{i\in[\overline{L}]}(\sqrt{\|\Sigma\|_{F}}\prod_{j=1}^{i}{\|\overline{\boldsymbol{W}}_{j}\|_{F}+\sum_{j=1}^{i}{\prod_{k=j+1}^{i-1}{\|\overline{\boldsymbol{W}}_{k}\|_{F}\|\overline{\boldsymbol{b}}_{j}\|_{2}}}}),\\ \sqrt{\|\Sigma\|_{F}})$. Under Assumption 1, there exists rank-$R$ or lower matrices $\bigtriangleup(\boldsymbol{W}_{l})_{l=1}^{L}\in\mathbb{R}^{D\times D}$ and bias vectors $(\hat{\boldsymbol{b}}_{l})_{l=1}^{L}\in\mathbb{R}^{D}$ for any input $\boldsymbol{x}\in\mathcal{X}$ with $\Sigma=\mathbb{E}\boldsymbol{xx}^{T}$, such that
(23) $\mathbb{E}\|f(\boldsymbol{x})-\overline{f}(\boldsymbol{x})\|_{2}\leqslant\xi\sum_{i=1}^{\overline{L}}{\max_{k\in[\overline{L}]}}(\|\overline{\boldsymbol{W}}_{k}\|_{F}+E_{k})^{\overline{L}-i}E_{i}.$
The detailed proofs of the aforementioned lemmas can be found in (Zeng and Lee, 2023), specifically within the proofs of Lemma 3, Theorem 3, and Theorem 5.
A.3. Expressive Power of Graph Neural Networks with LoRA
Assumption 2.
For a fixed rank $R\in[D]$, the weight matrices of the frozen model $(\boldsymbol{W}^{l})_{l=1}^{L}$ and matrices $(\prod_{l\in P_{i}}{\boldsymbol{W}^{l}})+LR_{r}(\overline{\boldsymbol{W}}^{i}-\prod_{l\in P_{i}}{\boldsymbol{W}^{l}})$ are non-singular for all $r\leqslant R(M-1)$ and $i\in[\overline{L}]$.
Theorem 4.
Under Assumption 2, if rank $R\geqslant\lceil\max_{i\in[\overline{L}]}rank(\overline{\boldsymbol{W}}^{i}-\prod_{l\in P_{i}}{\boldsymbol{W}^{l}})/M\rceil$, then there exists rank-$R$ or lower matrices $\bigtriangleup\boldsymbol{W}^{1},\cdots,\\ \bigtriangleup\boldsymbol{W}^{L}\in\mathbb{R}^{D\times D}$ and bias vectors $\hat{\boldsymbol{b}}^{1},\cdots,\hat{\boldsymbol{b}}^{L}\in\mathbb{R}^{D\times 1}$ such that the low-rank adapted model $g$ can exactly approximate the target model $\overline{g}$, i.e., $g(\boldsymbol{X})=\overline{g}(\boldsymbol{X})$, $\forall\boldsymbol{X}\in\mathcal{X}^{{}^{\prime}}$, where $\mathcal{X}^{{}^{\prime}}$ is the input space.
Proof.
First, let’s revisit the expression of the GNN: $\boldsymbol{H}^{l}=\mathrm{ReLU}(\boldsymbol{W}^{l}\boldsymbol{H}^{l-1}\boldsymbol{P}+\boldsymbol{B}^{l})$. We can observe that the GNN expression is akin to the FNN expression. However, there are two differences between the expressions of GNN and FNN. First, the input of the FNN is a single sample, while the input of the GNN includes all samples. Second, the GNN requires message passing at each layer: $\bar{\boldsymbol{H}}^{l}=\boldsymbol{H}^{l-1}\boldsymbol{P}$. We set $\boldsymbol{W}_{l}=\boldsymbol{W}^{l}$ for $l\in[L]$ and $\overline{\boldsymbol{W}}_{i}=\overline{\boldsymbol{W}}^{i}$ for $i\in[\overline{L}]$. Under Assumption 2, it is straightforward to observe that $(\boldsymbol{W}_{l})_{l=1}^{L}$ and $(\overline{\boldsymbol{W}}_{i})_{i=1}^{\overline{L}}$ satisfy Assumption 1. Despite the aforementioned differences between FNN and GNN, the proof process of Lemma 2 (Zeng and Lee, 2023) demonstrates that these differences do not affect the proof of the conclusion. Consequently, we can easily deduce that Theorem 4 holds. ∎
It is noteworthy that, based on Lemma 1, Assumption 2 holds in most cases. Revisiting the conditions for Theorem 4 to hold, if $R$ fails to meet the condition, can we offer an approximate upper bound on the difference between $g(\boldsymbol{X})$ and $\overline{g}(\boldsymbol{X})$? The answer is affirmative.
Theorem 5.
Define $E^{i}=\sigma_{RM+1}(\overline{\boldsymbol{W}}^{i}-\prod_{l\in P_{i}}{\boldsymbol{W}^{l}})$, and $\xi^{{}^{\prime}}=\max(\max_{i\in[\overline{L}]}(\mathbb{E}\|\boldsymbol{X}\|_{2}\prod_{j=1}^{i}{\|\overline{\boldsymbol{W}}^{j}}\|_{2}\|\boldsymbol{P}\|_{2}^{i}+\sum_{j=1}^{i}{\prod_{k=j+1}^{i-1}{\|\overline{\boldsymbol{W}}^{k}\|_{2}}}\\ \|\overline{\boldsymbol{B}}^{j}\|_{2}\|\boldsymbol{P}\|_{2}^{i-j-1}),\mathbb{E}\|\boldsymbol{X}\|_{2})$. Under Assumption 2, there exists rank-$R$ or lower matrices $\bigtriangleup(\boldsymbol{W}^{l})_{l=1}^{L}\in\mathbb{R}^{D\times D}$ and bias vectors $(\hat{\boldsymbol{b}}^{l})_{l=1}^{L}\in\mathbb{R}^{D\times 1}$ for any input $\boldsymbol{X}\in\mathcal{X}^{{}^{\prime}}$, such that
(24) $\mathbb{E}\|g(\boldsymbol{X})-\overline{g}(\boldsymbol{X})\|_{2}\leqslant\xi^{{}^{\prime}}\sum_{i=1}^{\overline{L}}{\max_{k\in[\overline{L}]}}(\|\overline{\boldsymbol{W}}^{k}\|_{2}+E^{k})^{\overline{L}-i}E^{i}\|\boldsymbol{P}\|_{2}^{\overline{L}-i+1}.$
Proof.
Similarly, we set $\boldsymbol{W}_{l}=\boldsymbol{W}^{l}$ for $l\in[L]$ and $\overline{\boldsymbol{W}}_{i}=\overline{\boldsymbol{W}}^{i}$ for $i\in[\overline{L}]$. Under Assumption 2, it is straightforward to observe that $(\boldsymbol{W}_{l})_{l=1}^{L}$ and $(\overline{\boldsymbol{W}}_{i})_{i=1}^{\overline{L}}$ satisfy Assumption 1. Following the proof of Lemma 3 (Zeng and Lee, 2023) and substituting the expression of GNN for the expression of FNN in the proof process of Lemma 3 (Zeng and Lee, 2023), it is easy to deduce that Theorem 5 holds. ∎
Appendix B Datasets
B.1. Dataset Statistics
Table 6. Statistics of datasets.
Dataset #Nodes #Edges #Features #Classes
PubMed 19,717 88,651 500 3
CiteSeer 3,327 9,228 3,703 6
Cora 2,708 10,556 1,433 7
Photo 7,650 238,163 745 8
Computer 13,752 491,722 767 10
Reddit 232,965 114,615,892 602 41
B.2. Dataset Descriptions
  • Citation networks. PubMed, CiteSeer, and Cora are three standard citation network benchmark datasets. In these datasets, the nodes correspond to research papers, and the edges represent the citations between papers. The node features are derived from the bag-of-words representation of the papers, while the node labels indicate the academic topics of the papers.
  • Co-purchase networks. The Photo and Computer datasets are segments of the Amazon co-purchase graph (McAuley et al., 2015). In these datasets, the nodes represent products, while the edges represent the frequent co-purchasing relationship between two products. The node features are derived from the bag-of-words representation of product reviews, and the labels indicate the categories of the products.
  • Reddit. The Reddit dataset is constructed from the Reddit online discussion forum. In this dataset, nodes represent posts, while edges indicate instances where the same user has commented on both connecting posts. The node features are derived from the GloVe CommonCrawl word representation of posts (Pennington et al., 2014), while labels indicate the community to which the post belongs.
Table 7. Comparison of experimental results in public and 10-shot settings. The notations ”-PM,” ”-CS,” ”-C,” ”-P,” and ”-Com” represent the pre-training datasets PubMed, CiteSeer, Cora, Photo, and Computer, respectively. The best experimental results are highlighted in bold, while the second-best results are underscored with a underline.
Method PubMed CiteSeer Cora Photo Computer
public 10-shot public 10-shot public 10-shot public 10-shot public 10-shot
GCN 78.66±0.56 73.28±0.93 70.50±0.75 64.52±1.40 82.00±0.97 71.88±1.26 92.17±0.75 85.39±1.57 86.66±1.20 71.97±1.16
GAT 78.30±0.43 74.96±0.67 70.94±1.08 64.14±2.12 80.58±1.50 72.04±0.61 92.91±0.22 86.38±0.65 86.80±0.71 74.82±2.36
COSTA 79.94±1.16 76.98±1.29 70.36±1.29 65.56±1.94 81.84±0.92 76.28±1.42 92.04±0.58 83.02±0.43 87.00±0.33 71.28±0.82
CCA-SSG 80.58±0.85 78.76±1.62 73.76±0.75 67.46±0.92 83.94±1.02 76.84±0.80 92.74±0.33 85.56±0.67 88.08±0.35 76.94±1.37
HomoGCL 81.04±0.05 79.42±0.04 71.38±0.04 65.40±0.00 82.40±0.00 75.70±0.00 92.43±0.08 84.30±0.35 87.75±0.29 76.85±0.41
GPPT 77.78±0.31 74.84±0.55 67.56±0.33 64.14±0.63 80.16±0.38 72.94±0.24 92.10±0.23 86.32±0.56 88.34±0.18 77.30±0.44
GRACE 79.52±0.16 75.86±0.11 70.34±0.21 67.70±0.00 82.28±0.04 76.40±0.00 92.32±0.31 86.16±0.02 85.54±0.30 74.39±0.03
GRACEt-PM 79.44±0.15 75.80±0.16 67.74±0.05 57.40±0.00 76.82±0.24 64.44±0.30 92.14±0.14 85.95±0.10 84.81±0.41 76.15±0.09
GRACEt-CS 76.58±0.04 72.46±0.39 70.50±0.24 67.70±0.00 79.04±0.05 71.70±0.00 92.46±0.38 86.81±0.02 85.08±0.62 76.05±0.01
GRACEt-C 73.00±0.00 66.78±0.04 67.10±0.12 58.14±0.13 82.32±0.04 76.40±0.00 92.23±0.21 86.12±0.03 84.46±0.24 75.33±0.30
GRACEt-P 71.10±0.00 57.70±0.62 58.00±0.00 49.38±0.04 72.40±0.00 57.88±0.04 92.25±0.34 86.20±0.02 84.24±0.21 74.57±0.01
GRACEt-Com 70.42±0.04 64.12±0.08 61.20±0.00 57.90±0.00 67.46±0.05 55.50±0.00 92.25±0.43 85.31±0.00 85.89±0.50 74.38±0.04
GTOT-PM 76.48±1.12 71.92±0.66 69.96±1.31 60.10±0.54 78.82±1.07 68.54±0.44 90.18±0.88 83.69±0.52 84.88±0.31 64.60±3.28
GTOT-CS 75.76±0.73 70.74±0.99 71.30±1.35 60.90±0.90 79.36±1.02 69.90±1.07 90.75±0.59 83.19±0.64 84.97±0.38 66.59±0.99
GTOT-C 75.66±0.80 70.72±0.30 68.98±1.01 60.98±0.89 79.36±0.78 69.84±0.86 90.24±1.33 84.21±0.86 85.30±0.11 66.46±0.59
GTOT-P 76.44±0.63 71.42±0.70 69.28±1.09 60.86±0.80 79.40±2.26 69.04±0.75 90.42±0.42 83.37±0.97 84.66±0.42 64.60±1.19
GTOT-Com 74.24±0.43 70.70±0.46 68.56±0.55 61.14±1.45 79.64±1.00 69.86±0.85 90.43±0.42 83.82±0.82 84.88±0.44 67.04±0.83
AdapterGNN-PM 76.44±0.97 72.78±0.72 62.76±1.42 58.64±0.44 75.54±1.46 63.82±1.44 92.39±0.32 88.24±0.39 88.00±0.18 75.54±0.65
AdapterGNN-CS 74.12±1.72 64.92±0.45 66.38±0.49 66.68±0.41 77.82±0.44 70.34±1.63 92.89±0.18 87.56±0.17 87.96±0.23 74.17±1.34
AdapterGNN-C 73.86±0.11 60.76±1.90 64.22±0.58 60.94±0.37 82.08±0.37 72.62±2.58 92.77±0.42 87.07±0.19 87.91±0.17 74.66±1.09
AdapterGNN-P 72.94±0.42 63.44±0.84 64.20±0.43 53.02±0.86 75.50±1.33 57.64±2.76 92.58±0.50 88.18±0.29 87.62±0.38 75.00±0.52
AdapterGNN-Com 72.50±0.62 58.94±2.89 63.64±0.68 58.74±0.33 74.12±0.73 56.42±2.50 92.66±0.45 88.20±0.77 87.63±0.49 72.64±1.96
GraphControl-PM 78.30±0.43 75.96±1.00 69.02±1.65 60.82±0.41 77.84±0.67 69.32±2.11 90.73±0.75 86.65±0.51 85.94±0.96 74.47±2.43
GraphControl-CS 75.98±0.66 72.56±0.63 70.80±0.97 68.56±0.98 77.54±1.24 74.04±0.79 90.15±0.67 87.44±0.29 86.36±0.26 74.03±1.01
GraphControl-C 74.52±0.88 66.00±0.66 66.20±0.94 58.70±0.56 77.14±1.72 76.44±0.31 90.52±0.48 86.57±0.70 85.99±0.51 73.17±1.50
GraphControl-P 74.58±1.99 58.94±0.69 59.12±1.34 53.36±1.92 73.46±1.73 65.72±0.44 90.67±0.50 86.23±0.59 86.11±0.50 71.86±2.38
GraphControl-Com 72.90±0.31 65.60±0.42 60.54±0.90 60.68±0.51 73.82±1.50 63.32±0.62 90.73±0.73 83.20±0.36 86.08±0.62 69.18±0.66
GraphLoRA-PM 80.86±0.39 78.06±0.59 74.20±0.47 74.62±0.57 82.42±0.40 78.08±0.3 93.00±0.36 88.34±0.51 87.70±0.63 76.54±0.39
GraphLoRA-CS 80.64±0.43 78.08±0.67 74.08±0.26 74.80±0.6 82.00±0.23 78.30±0.46 93.08±0.11 89.00±0.43 87.72±0.45 76.44±0.05
GraphLoRA-C 80.38±0.50 77.78±0.55 73.98±0.45 74.62±0.65 82.00±0.80 78.00±0.51 92.92±0.29 88.69±0.35 88.10±0.31 76.45±0.48
GraphLoRA-P 78.46±0.67 74.84±1.67 72.80±0.58 72.02±1.64 81.28±0.40 76.54±0.54 92.47±0.33 88.89±0.89 87.35±0.49 75.27±0.22
GraphLoRA-Com 79.02±0.77 74.44±1.36 72.72±0.35 72.12±1.29 80.84±1.05 76.44±0.9 92.42±0.22 89.07±0.22 87.28±0.31 75.35±1.26
Table 8. Comparison of experimental results in the 5-shot setting. The notations ”-PM,” ”-CS,” ”-C,” ”-P,” and ”-Com” represent the pre-training datasets PubMed, CiteSeer, Cora, Photo, and Computer, respectively. The best experimental results are highlighted in bold, while the second-best results are underscored with a underline.
Method PubMed CiteSeer Cora Photo Computer
GCN 70.00±0.67 61.92±2.12 67.78±0.61 86.10±0.65 68.79±2.36
GAT 69.98±0.34 62.42±1.43 68.56±1.41 86.86±0.71 73.23±1.38
COSTA 71.28±0.91 67.26±2.44 74.58±0.88 83.30±0.80 67.39±0.59
CCA-SSG 72.02±0.80 67.78±1.60 75.08±0.88 86.82±1.03 76.01±1.11
HomoGCL 70.30±0.00 67.50±0.10 75.44±0.00 83.73±0.18 77.64±0.48
GPPT 69.68±0.37 60.62±0.66 66.80±0.53 85.15±0.44 78.16±0.68
GRACE 69.38±0.08 64.08±0.13 72.82±0.44 87.53±0.01 71.64±0.13
GRACE-P 69.36±0.17 52.68±0.66 64.38±0.08 85.76±0.02 73.45±0.12
GRACE-CS 61.62±0.18 64.10±0.07 61.48±0.04 85.82±0.01 74.40±0.03
GRACE-C 62.24±0.17 54.44±0.09 72.74±0.05 85.45±0.00 73.05±0.56
GRACE-P 63.28±0.18 50.66±0.05 54.30±0.00 87.55±0.04 72.17±0.02
GRACE-Com 58.80±0.28 52.24±0.05 49.26±0.05 84.07±0.00 71.57±0.16
GTOT-PM 69.86±0.57 60.20±1.66 63.14±1.25 78.92±2.99 65.69±1.67
GTOT-CS 68.76±1.42 60.98±1.83 61.62±0.99 77.75±1.59 65.70±1.71
GTOT-C 69.04±0.96 61.52±1.03 61.63±0.89 79.64±1.40 66.53±1.12
GTOT-P 68.98±0.72 58.42±3.27 61.64±1.56 79.88±1.43 65.16±1.11
GTOT-Com 68.96±1.67 60.50±4.51 61.65±0.86 80.16±0.76 69.31±0.97
AdapterGNN-PM 69.44±0.59 54.94±0.72 60.04±2.09 87.03±0.41 74.21±0.55
AdapterGNN-CS 61.14±1.03 60.84±0.41 58.06±1.16 87.58±0.44 72.41±0.53
AdapterGNN-C 63.76±0.38 53.40±1.70 69.52±1.14 86.76±0.14 70.85±0.47
AdapterGNN-P 66.56±1.13 51.42±1.02 54.84±1.76 86.63±0.40 70.96±0.61
AdapterGNN-Com 60.36±0.89 52.86±0.49 51.54±2.14 86.93±0.18 69.95±0.80
GraphControl-PM 69.06±0.67 55.60±1.62 65.20±1.13 85.00±0.89 75.53±1.16
GraphControl-CS 63.72±0.77 64.38±0.35 63.62±1.55 84.64±1.08 74.42±1.06
GraphControl-C 67.14±1.06 55.06±1.33 73.24±1.20 84.72±0.94 73.82±0.72
GraphControl-P 65.54±0.48 53.30±1.20 56.36±0.85 87.36±1.09 72.32±1.02
GraphControl-Com 58.94±0.39 56.78±1.34 51.32±1.03 83.86±0.69 72.08±1.36
GraphLoRA-PM 72.56±0.96 72.52±2.50 77.16±0.62 86.05±0.19 75.57±0.64
GraphLoRA-CS 73.54±1.66 71.94±2.46 76.98±0.32 86.51±0.66 75.90±0.45
GraphLoRA-C 71.34±0.40 71.78±2.11 77.00±0.25 86.24±0.13 75.47±0.35
GraphLoRA-P 69.02±0.72 71.32±1.91 74.38±1.31 86.57±0.71 74.63±1.38
GraphLoRA-Com 69.30±1.68 70.72±0.64 73.70±2.00 86.34±0.81 73.81±1.24
Table 9. Ablation experiment results in the 5-shot setting. The best experimental results are highlighted in bold.
Variants PubMed CiteSeer Cora Photo Computer
w/o mmd 72.12±0.73 71.18±2.38 76.02±0.85 85.83±0.18 75.24±0.20
w/o smmd 69.70±0.14 69.00±5.13 74.96±1.72 85.26±0.35 74.69±0.82
w/o cl 71.28±0.79 70.76±4.99 75.90±0.49 86.07±0.50 75.05±0.97
w/o str 72.66±0.78 66.08±0.58 65.84±1.30 85.73±0.46 74.63±1.18
w/o lrd 71.06±0.88 70.84±1.57 74.68±1.24 86.06±0.17 74.44±0.36
w/o nfa 69.82±0.26 33.94±0.21 69.52±0.28 77.14±0.99 73.53±0.35
w/o sktl 71.14±0.61 64.40±4.87 75.78±0.30 85.73±0.07 74.91±0.57
GraphLoRA 72.56±0.96 72.52±2.50 77.16±0.62 86.05±0.19 75.57±0.64
Appendix C Baselines
  • GCN (Kipf and Welling, 2017): GCN is a foundational graph neural network that effectively propagates information within the graph structure by capturing the relationships among nodes and their neighboring nodes.
  • GAT (Velickovic et al., 2018): GAT is another classic graph neural network. In contrast to GCN, GAT introduces attention mechanisms, allowing each node to dynamically adjust weights based on the importance of its neighboring nodes during the representation update process.
  • GRACE (Zhu et al., 2020): GRACE adopts the SimCLR framework (Chen et al., 2020a) and incorporates two strategies to augment the source graph. It aims to maximize the mutual information between two views by enhancing agreement at the node level.
  • COSTA (Zhang et al., 2022b): COSTA introduces a feature augmentation framework to perform augmentations on the hidden features, mitigating the issue of highly biased node embeddings obtained from graph enhancement. Moreover, it accelerates the speed of graph contrastive learning.
  • CCA-SSG (Zhang et al., 2021): CCA-SSG proposes an innovative feature-level optimization objective based on Canonical Correlation Analysis for graph contrastive learning, presenting a conceptually simple yet effective model.
  • HomoGCL (Li et al., 2023b): HomoGCL enhances graph contrastive learning by leveraging the homophily of the graph. It directly utilizes the homophily of the graph by estimating the probability of neighboring nodes being positive samples via a Gaussian Mixture Model.
  • GPPT (Sun et al., 2022): GPPT is a graph prompt learning method that introduces a novel paradigm for graph neural network transfer learning known as ”pre-train, prompt, fine-tune”, designed specifically for cross-task transfer learning.
  • GRACEt: GRACEt is a variant of GRACE that involves pre-training on the source graph using the GRACE method and fine-tuning on the target graph.
  • GTOT (Zhang et al., 2022a): GTOT is an optimal transport-based fine-tuning method. It formulates graph local knowledge transfer as an optimal transport problem, preserving the local information of the fine-tuned network from pre-trained models.
  • AdapterGNN (Li et al., 2024): AdapterGNN is a parameter-efficient fine-tuning method, which freezes the pre-trained network and introduces adapters to it.
  • GraphControl (Zhu et al., 2023): GraphControl is a recent research endeavor in the field of graph neural network transfer learning. Drawing inspiration from ControlNet (Zhang et al., 2023), it incorporates its core concepts to enhance transfer learning in graph neural networks.
Appendix D Experimental settings
For supervised methods, we utilize labeled data from the target graph to train the GNN in a supervised manner. On the other hand, for graph contrastive learning methods and graph prompt learning methods, we first pre-train the GNN on the target graph in an unsupervised manner. Subsequently, we freeze the model parameters and fine-tune either a linear classifier or a graph prompt using the labels from the target graph. For transfer learning methods, we adopt a similar approach by pre-training the GNN on the source graph in an unsupervised manner. We then transfer the pre-trained model to the target graph and fine-tune a linear classifier using the labels from the target graph.
For all baselines, we utilize publicly available code provided by the authors. We employ the Adam (Kingma and Ba, 2015) optimizer for optimization, and the hyperparameters are tuned as suggested by the authors.
Appendix E Additional Experimental Results
E.1. Performance Comparison in the Public Setting and the 10-shot Setting
The comprehensive results of the comparison experiment in the public setting and the 10-shot setting are shown in Table 7.
E.2. Performance Comparison in the 5-shot Setting
The results of the comparison experiment in the 5-shot setting are shown in Table 8.
E.3. Ablation Studies in the 5-shot Setting
The results of the ablation experiment in the 5-shot setting are shown in Table 9.